OpenStudy (anonymous):

Factoring by grouping 2x^2+4-2

OpenStudy (anonymous):

Is it \[2x ^{2}+4x-2\] or just \[2x^{2}+4-2\]

OpenStudy (anonymous):

sorry, its the first one

OpenStudy (anonymous):

sorry, i made you wait so long but this can't be factored by grouping..

OpenStudy (anonymous):

so would that make it 2 complex solutions or 2 irrational solutions

OpenStudy (campbell_st):

I still think there is an error as the expression can't be grouped in pairs to factor... the only common factor is 2 then its 2(x^2 + 2x -1)

OpenStudy (anonymous):

Well the section im in is factoring by grouping but it says to predict the solutions a. two rational solutions b.one rational solution c. two irrational solutions d. two complex solutions

OpenStudy (campbell_st):

use the discriminant to determine the nature of the roots

OpenStudy (anonymous):

-b-4ac? 2x^2+4x-2 (4)^2 -4(2)(-2) 16-8(-2) 16-16 0 right?

OpenStudy (campbell_st):

well almost 16 - 4(2)(-2) = 16 + 16 = 32 so the roots are real and irrational...

OpenStudy (anonymous):

you saved my life!